OpenStudy (lgbasallote):

Lgbariddle >:))) take 9 from 6 take 6 from 7 take 10 from 20 and the result is 6 \(\large \mathbb{\text{how is this so???}}\)

OpenStudy (lgbasallote):

Hint: this needs an out-of-the-box thinking >:))

OpenStudy (lgbasallote):

\(\Huge \mathbb{\text{^spam}}\)

OpenStudy (saifoo.khan):

Lol, stop using \(\LaTeX\)

OpenStudy (mimi_x3):

\[\large -3 +-1+10 = 6 \] lol, just a guess..but got the answer answer..

OpenStudy (mimi_x3):

correct answer*

OpenStudy (lgbasallote):

lol no @Mimi_x3 :P how did you get -1...the way you went it was supposed 7-6 = 1 :P

OpenStudy (dumbcow):

technically then it should say "7 from 6" then for that logic to work

OpenStudy (lgbasallote):

@saifoo.khan you can't stop the \(\mathbf{\LaTeX}\)

OpenStudy (anonymous):

3 X 2 = 6

OpenStudy (lgbasallote):

nice job :DDD \(\Huge \checkmark\)

OpenStudy (mimi_x3):

lol, what kind of riddle is this..how stupid..

OpenStudy (apoorvk):

To be honest, I had come across a similar riddle long time back, (the nos. and the results were different). I just remembered the method somehow, and it clicked at the right time, so, no points for me :| Please don't gimme a medal. :|

OpenStudy (lgbasallote):

lol mimi =)))))

OpenStudy (lgbasallote):

ill think of a harder one for you @Mimi_x3 :P

OpenStudy (mimi_x3):

that relates to mathematics please..that was stupid..

OpenStudy (dumbcow):

good ol roman numerals, should have thought of that

OpenStudy (lgbasallote):

ohkay! :DD

OpenStudy (saifoo.khan):

Outside the boc